OpenStudy (anonymous):

Carleigh makes $10 more then Sheryl (s) does per hour. Carleigh makes $22 an hour. What an equation to represent this situation

OpenStudy (anonymous):

22-10=s maybe not for sure

OpenStudy (anonymous):

the choices are s+10=22 and 10+22=s

OpenStudy (anonymous):

s+10=22